OpenStudy (anonymous):

prove that cosx / 1+sinx = 1-sinx / cosx

myininaya (myininaya):

\[\frac{\cos(x)}{1+\sin(x)} \cdot \frac{1-\sin(x)}{1-\sin(x)}=\frac{\cos(x)(1-\sin(x))}{1-\sin^2(x)}\]

myininaya (myininaya):

1-sin^2(x)=cos^2(x)

myininaya (myininaya):

so something cancels right?

myininaya (myininaya):

\[\frac{\cos(x)(1-\sin(x))}{\cos^2(x)}=\frac{1-\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

thank you !!!